OpenStudy (anonymous):

for what positive value of a does a-2(sqrta+7) equals -4 ??

OpenStudy (anonymous):

\[a-2\sqrt{a+7} equals -4\]

OpenStudy (unklerhaukus):

what number would you like to take the square root of?

OpenStudy (anonymous):

sorry.. i dont get you .

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

it is easy to take the square root of a square number like 1, 4, ... √1=√1^2=1 √4=√2^2=2 .. ..

OpenStudy (anonymous):

so we can plug in any sq. no. ???

OpenStudy (anonymous):

got it :) ty

OpenStudy (unklerhaukus):

other numbers wont solve the equation